OpenStudy (anonymous):

evaluate 19-5/ (1/3 + 1/2)

OpenStudy (anonymous):

1/3 + ½=2/6+3/6=5/6 5/(5/6)=5/1*6/5=6 19-6=13

OpenStudy (anonymous):

Please show the different answer.

OpenStudy (anonymous):

1) Order of operation depends on the way how parentheses are set. I assumed the atlasatlas wrote it as in a question. 2) Even if numerator is (19-5), final answer will be 84/5 = 16.8 because 14/(5/6)=(14/1)/(5/6)= (14/1)*(6/5)= 84/5 = 16.8

OpenStudy (anonymous):

I agree with answer = 16.8, exactly how Marina set it out in her last entry.